OpenStudy (anonymous):

inverse of f(x)=-2/5x-3

OpenStudy (anonymous):

y=-2/5x-3 y-3=-2/5x (-5/2)(y-3)=x -- now change x with y (-5/2)(x-3)=y f inverse of x = (-5/2)(x-3) = (-5/2)x + (-15/2)

OpenStudy (anonymous):

oops, + 15/2 not +(-15/2) sorry